An organ pipe is made to play a low note at 27.5 Hz, the same as the lowest note on a piano.

Part A

Assuming a sound speed of 343 m/s, what length open-open pipe is needed?

Express your answer with the appropriate units.

Part B

What length open-closed pipe would suffice?

Express your answer with the appropriate units.
Answer #1
Answer #2

λ = v / f = 343 m/s / 27.5 s = 12.5 m

A) open ended, L = λ/2, so
L = 12.5 m / 2 = 6.24 m

B) closed ended, L = λ / 4, so
L = 12.5 m / 4 = 3.12 m

answered by: Dr. R